# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
The recent discovery of superconductivity with a transition temperature $T_{c}$ over 40 K in La3Ni2O7 and (La,Pr)3Ni2O7 thin films at ambient pressure marks an important step in the field of nickelate superconductors. Here, we perform a renormalized mean-field theory study of the superconductivity in $\mathrm{La_{3}Ni_{2}O_{7}}$ thin films, using a bilayer two-orbital $t-J$ model. Our result reveals an $s_{\pm}$-wave pairing symmetry driven by the strong interlayer superexchange coupling of $d_{z^{2}}$ orbital, resembling the pressurized bulk case. Also, we roughly reproduce the experimentally reported nodeless shape of the superconducting gap at the $\beta$ pocket and the superconducting $T_{c}$. To gain insight into the orbital feature of superconductivity, we explore the projection of different pairing bonds on Fermi surface. We find that the nodeless gap at $\beta$ pocket is related to the interlayer pairing within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals, in which the latter is triggered by the former through hybridization, and both hold the same sign.
@@
In this paper, using a bilayer $t-J$ model that captures the key ingredient of LNO thin films, we perform a renomalized mean-field theory (RMFT) [7] calculation of the superconductivity. Our result reveals an $s_{\pm}$-wave pairing and reproduces the nodeless shape of the superconducting gap at $\beta$ pocket as reported in Ref. [60]. Apparently, the presence of a finite gap along BZ diagonal direction has indicated a direct $s_{\pm}$-wave pairing on $d_{x^{2}-y^{2}}$ orbital. In general, the gap structure on Fermi surface is sensitive to the relative strengths of pairing configurations in real space, and our result suggests that such a gap shape is a joint effort of the interlayer pairing bonds within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals.
@@
In this model, the basis is $\Psi=(d_{z_{1}},d_{x_{1}},d_{z_{2}},d_{x_{2}})^{T}$ that contains $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals for the first and second layers, as shown in Fig. 1(b). $\mathcal{H}_{t}$ is the tight-binding Hamiltonian adopted from DFT simulation of the double-stacked LNO thin films [57]. Here, $ij,\mu\nu,\sigma$ denote the lattice, orbital, and spin indices, respectively. The band structure is shown in Fig. 1(a), in which a small electron doping level with chemical potential $\mu=-0.05$ is used in our calculation to better agree with the experiments, which gives the elctron filling $n_{z}$=0.8, $n_{x}$=0.58. For the Heisenberg term $\mathcal{H}_{J}$, $J_{\bot}$ is the interlayer superexchange coupling of $d_{z^{2}}$ orbitals, and $J_{||},J_{xz}$ is the inplane nearest neighbor superexchange coupling of $d_{x^{2}-y^{2}}$ orbital, and superexchange coupling between $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$, respectively. Also, the Hund’s coupling $J_{H}$ is included for generality. To quantitatively estimate $J_{\bot},J_{||},J_{xz}$, we perform the exact diagonalization (ED) calculation. Taking $J_{\bot}$ as an example, ED is performed for a 5-site chain containing two $d_{z^{2}}$ and three $p_{z}$ orbitals along the $z$ axis, where the $p-d$ hoppings are also obtained from the DFT calculation [57]. Under the atomic subspace ($N$=8, $S_{z}$=0) that corresponds to half-filling of $d_{z^{2}}$ orbital, $J_{\bot}$ is determined as the energy difference of the lowest singlet and triplet states [8]. We finally obtain $J_{\bot}$=0.135 eV, $J_{||}$=0.084 eV, $J_{xz}$=0.03 eV. Apart from this, a realistic $J_{H}$=1 eV is also used in this study.
@@
Pairing symmetry and superconducting gap.-- Figure. 2(a) presents the superconducting gap matrix $[g_{t}\Delta]_{\mu\nu}$ projected onto FS, which demonstrates that $\beta,\gamma$ pockets are nodeless and of opposite sign, while $\alpha$ pocket shows a sign flip. To better illusstrate the gap features, Fig. 2(b-d) show the absolute value $g_{t}|\Delta|$ as a function of angle $\theta$ for $\alpha,\beta$ and $\gamma$ pockets, respectively. For $\alpha$ pocket in Fig. 2(b), the gap demonstrates a node between $\theta=45^{\circ}\sim 90^{\circ}$. For $\beta$ pocket in Fig. 2(c), the gap shows a parabolic shape centering at $\theta=45^{\circ}$, which gets slightly decrease when approaching the corner. Note that the gap shape as well as its magnitude around the BZ diagonal direction from our result are generally coordinated with the ARPES result [60], which suggests that our RMFT captures the key ingredient in the LNO thin films. For $\gamma$ pocket in Fig. 2(d), we find the gap magnitude is around 5 times larger than that in Fig. 2(b), and a minimum is also identified at the BZ diagonal direction. Apparently, such large gap opening at $\gamma$ pocket should be attributed to the strong interlayer superexchange coupling $J_{\bot}$, which is also consistent with previous calculations for bulk LNO [18, 16, 25, 7]. Overall, the pairing structure in Fig. 2 should be described as the $s_{\pm}$-wave, which can also be seen from our result that besides the formation of the large interlayer pairng bond for $d_{z^{2}}$ orbital ($\Delta_{\bot}^{z}$), the inplane pairing bonds have the same magnitude when along $x,y$ directions for both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals ($\Delta_{||}^{x},\Delta_{||}^{z})$. Note that, in our $t-J$ model only the pairing bonds $\Delta_{\bot}^{z},\Delta_{||}^{x},\Delta_{||}^{xz}$ are associated with the applied $J_{\bot},J_{||},J_{xz}$, respectively, while other terms like the interlayer pairing of $d_{x^{2}-y^{2}}$ orbital ($\Delta_{\bot}^{x}$) do not have the corresponding $J$ and should not appear in the calculation. However, they can still gain finite values from the mean-field wave function as $\Delta_{ij}^{\mu\nu}=\langle n{\rm k}|c_{i\mu\uparrow}^{\dagger}c^{\dagger}_{j\nu\downarrow}|n{\rm k}\rangle$, which reflects the pairing tendency under a global coherent state. For this reason, in Fig. 2 we have incorporated all the relevant $\Delta$, and in the following we will illustrate the role of different $\Delta$ on FS.
@@
Orbital-resolved superconducting gap.-- Figure. 3 presents the orbital-resolved superconducting gap for (a) $\Delta_{\bot}^{z}$, (b) $\Delta_{||}^{z}$, (c) $\Delta_{\bot}^{x}$, and (d) $\Delta_{||}^{x}$, respectively. For $\Delta_{\bot}^{z}$ in Fig. 3(a), the gap is positive for the bonding branch of $\alpha,\gamma$ pockets, and is negative for the anti-bonding branch of $\beta$ pocket. For $\Delta_{||}^{z}$ in Fig. 3(b), the gap structure clearly demonstrates the sign-flip $s_{\pm}$-wave with the node appearing at the reduced BZ boundary. Also, another node at the BZ diagonal direction in both Fig. 3(a,b) is identified due to the vanishing of hybridization. For $\Delta_{\bot}^{x},\Delta_{||}^{x}$ in Fig. 3(c,d), similar behavior is observed as compared to Fig. 3(a,b), but with a much smaller magnitude. Notably, the sign in Fig. 3(d) is opposite to that in Fig. 3(b), namely, $sgn(\Delta_{||}^{x})=-sgn(\Delta_{||}^{z})$. This feature is also seen in a slave-boson mean-field study [31], and can be explained as a result of the competition between $J_{\bot}$ preferred $s_{\pm}$-wave and $J_{||}$ preferred $d$-wave pairings [64]. Finally, combining Fig. 3(a,c), it is evident that the experimentally observed nodeless gap at $\beta$ pocket is a joint effort of the interlayer pairing within both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals. More insights on gap structures with the real space pairing configurations are provided in supplementary [64].
@@
Transition temperature.--In Fig. 4, we show the superconducting gap as a function of temperature $T$, which reveals two dominated pairing bonds from $d_{z^{2}}$ orbitals (red), in line with Fig. 3. As increasing $T$, these gaps decrease in a mean-field manner and simultaneously drop to zero at $T_{c}$$\approx$60 K. Such value is comparable to the experimental reported $T_{c}$=40$\sim$60 K in the thin films [54, 55, 50]. Given that our previous RMFT study also obtains an experimental comparable $T_{c}\approx 80$ K for the pressurized bulk LNO [7], it is reasonable to believe that the decrease of $T_{c}$ in the thin films should originate from the decrease of $J_{\bot}$. This speculation draws us to the difference in the electronic structure, especially in the mainly concerned interlayer hopping $t_{\bot}^{z}$ for $d_{z^{2}}$ orbital. Intriguingly, we find that the $J_{\bot}$=0.135 eV in this study is $\sim 25\%$ smaller than that of $J_{\bot}$=0.18 eV for the pressurized bulk [8] as estimated from ED, which is coordinated with a simple estimation using $J_{\bot}\propto(t_{\bot}^{z})^{2}$ [4, 57]. Recently, a strain-tuning experiment in LNO thin films has reported an increase of $T_{c}$ from 10 K in the tensile-strained $\mathrm{SrTiO_{3}}$ substrate to 60 K in the compressively strained $\mathrm{LaAlO_{3}}$ substrate under 20 GPa of pressure [50]. This indicates a positive dependence of $T_{c}$ on the lattice ratio $c/a$ and, at first glance, violates the above understanding. However, we believe $J_{\bot}$ can only determine the upper bound of $T_{c}$, while other factors like the hybridization, density wave, and interlayer Josephson coupling can also have intricate impact, as also pointed out in Ref. [65, 66, 67, 68, 69, 50, 70].
@@
Discussion.-- Our RMFT study of the superconductivity in LNO thin films has predicted an $s_{\pm}$-wave pairing symmetry. The result largely resembles that of the pressurized bulk LNO [18, 16, 25, 7] giving the similarity in the electronic structure. Mathematically, our application of Gutzwiller projection  [62, 18] on a multi-orbital $t-J$ model emposes a constraint on the renormalization condition so as to bring a simplification to have the quadratic form of the Hamiltonian, hence can be solved self-consistently like the standard Bardeen-Cooper-Schrieffer (BCS) theory. For this reason, it is expected that the obtained $s_{\pm}$-wave mainly reflects pairing preference for the FS toplogy in LNO [71]. Nevertheless, we find that the $s_{\pm}$-wave is quite robust under various approaches from weak to strong couplings, which convinces us that the fluctuation of correlated atomic states might not significantly alert such picture. In this picture, the $\gamma$ pocket plays a central role. Experimentally, there are two relevant ARPES measurements in which the $\gamma$ pocket shows visible profile for that performed on the normal state [56] but invisible for that on the superconducting state [51]. From our result, this contradiction might be due to the opening of a large gap that exceeds the detected energy range, as shown Fig. 2(d). Further experimental evidences are expected on this issue.
@@
Conclusion-- To summarize, we perform a RMFT study of the superconductivity in the LNO thin films using a bilayer $t-J$ model. Our result reveals an $s_{\pm}$-wave pairing symmetry driven by the interlayer superexchange coupling, resembling the bulk LNO case. Our result also roughly reproduces the experimentally reported nodeless shape of the superconducting gap at $\beta$ pocket  [60] as well as the superconducting $T_{c}$. In addition, we systematically analyse the relation between orbital-resolved pairing bonds and the gap structure on FS. The result shows that the nodeless gap of $\beta$ pocket is related to the interlayer pairing within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals, in which the latter is triggered by the former through hybridization. Both of them hold the same sign.
@@
We thank the useful discussion with Wéi Wú. This project was supported by NSFC-12494591, NSFC-92165204, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Guangdong Fundamental Research Center for Magnetoelectric Physics (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).

### v2

@@ Abstract
The recent discovery of superconductivity with a transition temperature $T_{c}$ over 40 K in La3Ni2O7 and (La,Pr)3Ni2O7 thin films at ambient pressure marks an important step in the field of nickelate superconductors. Here, we perform a renormalized mean-field theory study of the superconductivity in $\mathrm{La_{3}Ni_{2}O_{7}}$ thin films, using a bilayer two-orbital $t-J$ model. Our result reveals an $s_{\pm}$-wave pairing symmetry driven by the strong interlayer superexchange coupling of $d_{z^{2}}$ orbital, resembling the pressurized bulk case. Also, we roughly reproduce the experimentally observed nodeless shape of the superconducting gap at the $\beta$ pocket and the superconducting $T_{c}$. In addition, by analysing the orbital-resolved pairing configurations and their projections onto Fermi surface, we find that the nodeless feature of $\beta$ pocket is related to the interlayer pairing within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals. Moreover, we identify a formation of the inplane inter-orbital $d$-wave pairing between $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals, which can even enhance the dominated interlayer $s_{\pm}$-wave.
@@
In this paper, using a bilayer $t-J$ model that captures the key ingredient of LNO thin films, we perform a renomalized mean-field theory (RMFT) [7] calculation of the superconductivity. Our result reveals an $s_{\pm}$-wave pairing and reproduces the experimentally observed nodeless shape of the superconducting gap at $\beta$ pocket in Ref. [60]. To gain insight, we also analyse the orbital-resolved pairing bonds and their projections onto Fermi surface.
@@
In this model, the basis is $\Psi=(d_{z_{1}},d_{x_{1}},d_{z_{2}},d_{x_{2}})^{T}$ that contains $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals for the first and second layers, as shown in Fig. 1(b). $\mathcal{H}_{t}$ is the tight-binding Hamiltonian adopted from DFT simulation for the double-stacked LNO thin films [57]. Here, $ij,\mu\nu,\sigma$ denote the lattice, orbital, and spin indices, respectively. The band structure is shown in Fig. 1(a), in which a small electron doping level with chemical potential $\mu=-0.05$ is used in our calculation to better agree with the experiments, which gives the electron filling $n_{z}$=0.8, $n_{x}$=0.58. For the Heisenberg term $\mathcal{H}_{J}$, $J_{\bot}$ is the interlayer superexchange coupling of $d_{z^{2}}$ orbitals, and $J_{||},J_{xz}$ is the inplane nearest neighbor superexchange coupling of $d_{x^{2}-y^{2}}$ orbital, and superexchange coupling between $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$, respectively. Also, the Hund’s coupling $J_{H}$ is included for generality. To quantitatively estimate $J_{\bot},J_{||},J_{xz}$, we perform the exact diagonalization (ED) calculation. Taking $J_{\bot}$ as an example, ED is performed for a 5-site chain containing two $d_{z^{2}}$ and three $p_{z}$ orbitals along the $z$ axis, where the $p-d$ hoppings are also obtained from the DFT calculation [57]. Under the atomic subspace ($N$=8, $S_{z}$=0) that corresponds to half-filling of $d_{z^{2}}$ orbital, $J_{\bot}$ is determined as the energy difference of the lowest singlet and triplet states [8]. We finally obtain $J_{\bot}$=0.135 eV, $J_{||}$=0.084 eV, $J_{xz}$=0.03 eV. Apart from this, a realistic $J_{H}$=1 eV is also used in this study.
@@
Pairing symmetry and superconducting gap.-- Figure. 2(a) presents the superconducting gap matrix $[g_{t}\Delta]_{\mu\nu}$ projected onto FS, which demonstrates that $\beta,\gamma$ pockets are nodeless and of opposite sign, while $\alpha$ pocket shows a sign flip. To better illusstrate the gap features, Fig. 2(b-d) show the absolute value $g_{t}|\Delta|$ as a function of angle $\theta$ for $\alpha,\beta$ and $\gamma$ pockets, respectively. For $\alpha$ pocket in Fig. 2(b), the gap demonstrates a node between $\theta=45^{\circ}\sim 90^{\circ}$. For $\beta$ pocket in Fig. 2(c), the gap shows a parabolic shape centering at $\theta=45^{\circ}$, which gets slightly decrease when approaching the corner. Note that the gap shape as well as its magnitude around the BZ diagonal direction from our result are generally coordinated with the ARPES result [60], which suggests that our RMFT captures the key ingredient in the LNO thin films. For $\gamma$ pocket in Fig. 2(d), we find the gap magnitude is around 5 times larger than that in Fig. 2(b), and a minimum is also identified at the BZ diagonal direction. Apparently, such large gap opening at $\gamma$ pocket should be attributed to the strong interlayer superexchange coupling $J_{\bot}$, which is also consistent with previous calculations for bulk LNO [18, 16, 25, 7]. Overall, the pairing structure in Fig. 2 should be described as the $s_{\pm}$-wave, which can also be seen from our result that besides the formation of the large interlayer pairng bond for $d_{z^{2}}$ orbital ($\Delta_{\bot}^{z}$), the inplane pairing bonds have the same magnitude when along $x,y$ directions for both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals ($\Delta_{||}^{x},\Delta_{||}^{z})$. Note that, in our $t-J$ model only the pairing bonds $\Delta_{\bot}^{z},\Delta_{||}^{x},\Delta_{||}^{xz}$ are associated with the applied $J_{\bot},J_{||},J_{xz}$, respectively, while other terms like the interlayer pairing of $d_{x^{2}-y^{2}}$ orbital ($\Delta_{\bot}^{x}$) do not have the corresponding $J$ and should not appear in the calculation. However, they can still gain finite values from the mean-field wave function as $\Delta_{ij}^{\mu\nu}=\langle n{\rm k}|c_{i\mu\uparrow}^{\dagger}c^{\dagger}_{j\nu\downarrow}|n{\rm k}\rangle$, which reflects the pairing tendency towards a global coherent state. For this reason, in Fig. 2 we have incorporated all the relevant $\Delta$, and in the following we will illustrate the role of different $\Delta$ on FS.
@@
Orbital-resolved superconducting gap.-- Figure. 3 presents the orbital-resolved superconducting gap for (a) $\Delta_{\bot}^{z}$, (b) $\Delta_{||}^{z}$, (c) $\Delta_{\bot}^{x}$, and (d) $\Delta_{||}^{x}$, respectively. For $\Delta_{\bot}^{z}$ in Fig. 3(a), the gap is positive for the bonding branch of $\alpha,\gamma$ pockets, and is negative for the anti-bonding branch of $\beta$ pocket. For $\Delta_{||}^{z}$ in Fig. 3(b), the gap structure clearly demonstrates the sign-flip $s_{\pm}$-wave with the node appearing at the extended BZ boundary (diamond shape). Also, another node at the BZ diagonal direction in both Fig. 3(a,b) is identified due to the vanishing of hybridization. For $\Delta_{\bot}^{x},\Delta_{||}^{x}$ in Fig. 3(c,d), similar behavior is observed as compared to Fig. 3(a,b), but with a much smaller magnitude. It is worth noting that the phases in Fig. 3(a-d) all can be understood in terms of maximizing the total gap magnitude. First, Fig. 3(c) will choose the same sign with that in Fig. 3(a) ($sgn\Delta_{\bot}^{z}$=$sgn\Delta_{\bot}^{x}$), which is spatially permissible. This also explains the experimentally observed nodeless shape of $\beta$ pocket as a joint effort of interlayer pairing within both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals. Second, due to the inplane $\cos k_{x}+\cos k_{y}$ form factor in Fig. 3(b,d), cancelation of the gap in certain region of FS is unavoidable, which inherently reflects the “pairing frustration”. In this case $\Delta_{||}^{z}$ will choose to maximize the gap for $\gamma$ pocket at the price of that for $\alpha$ pocket, while $\Delta_{||}^{x}$ will no doubt choose to maximize the gap for $\alpha$ pocket, with slight cancelation at $\beta$ pocket. These result in: $sgn\Delta_{\bot}^{z}=sgn\Delta_{\bot}^{x}=sgn\Delta_{||}^{x}=-sgn\Delta_{||}^{z}$. Similar features are also seen in Ref. [18, 31] for bulk LNO. Finally, concerning the inplane inter-orbital pairing $\Delta_{||}^{xz}$, surprisedly, we find it actually forms a $d$-wave and can even enhance the dominated $\Delta_{\bot}^{z}$. This is a unique feature for LNO $I4/mmm$ lattice where the two distinct pairings can coexist within $A_{1g}$ channel, see supplementary [64] and Ref. [65] for details.
@@
Transition temperature.--In Fig. 4, we show the superconducting gap as a function of temperature $T$, which reveals two dominated pairing bonds from $d_{z^{2}}$ orbitals (red), in line with Fig. 3. As increasing $T$, these gaps decrease in a mean-field manner and simultaneously drop to zero at $T_{c}$$\approx$60 K. Such value is comparable to the experimental reported $T_{c}$=40$\sim$60 K in the thin films [54, 55, 50]. Given that our previous RMFT study also obtains an experimental comparable $T_{c}\approx 80$ K for the pressurized bulk LNO [7], it is reasonable to believe that the decrease of $T_{c}$ in the thin films should originate from the decrease of $J_{\bot}$. This speculation draws us to the difference in the electronic structure, especially in the mainly concerned interlayer hopping $t_{\bot}^{z}$ for $d_{z^{2}}$ orbital. Notably, we find that the $J_{\bot}$=0.135 eV in this study is $\sim 25\%$ smaller than that of $J_{\bot}$=0.18 eV for the pressurized bulk [8] as estimated from ED, which is coordinated with a simple estimation using $J_{\bot}\propto(t_{\bot}^{z})^{2}$ [4, 57]. Recently, a strain-tuning experiment in LNO thin films has reported an increase of $T_{c}$ from 10 K in the tensile-strained $\mathrm{SrTiO_{3}}$ substrate to 60 K in the compressively strained $\mathrm{LaAlO_{3}}$ substrate under 20 GPa of pressure [50]. This indicates a positive dependence of $T_{c}$ on the lattice ratio $c/a$ and, at first glance, violates the above understanding. However, we believe $J_{\bot}$ can only determine the upper bound of $T_{c}$, while other factors like the hybridization, density wave, and interlayer Josephson coupling can also have intricate impact, as also pointed out in Ref. [66, 67, 68, 69, 70, 50, 71].
@@
Discussion.-- Our RMFT study of the superconductivity in LNO thin films has predicted an $s_{\pm}$-wave pairing symmetry. The result largely resembles that of the pressurized bulk LNO [18, 16, 25, 7] giving the similarity in the electronic structure. Mathematically, our application of Gutzwiller projection  [62, 18] on a multi-orbital $t-J$ model imposes a constraint on the renormalization condition so as to bring a simplification to have the quadratic form of the Hamiltonian, hence can be solved self-consistently like the standard Bardeen-Cooper-Schrieffer (BCS) theory. For this reason, it is expected that the obtained $s_{\pm}$-wave mainly reflects pairing preference for the LNO FS topology in the presence of $J_{\bot}$ [72]. Nevertheless, we find that the $s_{\pm}$-wave is quite robust under various approaches from weak to strong couplings, which convinces us that the fluctuation of correlated atomic states might not significantly alert such scenario. In this scenario, the $\gamma$ pocket plays a central role. Experimentally, there are two relevant ARPES measurements in which the $\gamma$ pocket shows visible profile for that performed in the normal state [56] but invisible for that in the superconducting state [51]. From our result, this contradiction might be due to the opening of a large gap that exceeds the detected energy range, as shown Fig. 2(d). More experimental evidences are expected on this issue. Regarding the role of $J_{H}$, no significant impact is observed from our approach.
@@
Conclusion-- To summarize, we perform a RMFT study of the superconductivity in the LNO thin films using a bilayer $t-J$ model. Our result reveals an $s_{\pm}$-wave pairing symmetry driven by the interlayer superexchange coupling, resembling the bulk LNO case. Our result also roughly reproduces the experimentally reported nodeless shape of the superconducting gap at $\beta$ pocket  [60] as well as the superconducting $T_{c}$. In addition, we systematically analyse the relation between orbital-resolved pairing bonds and the gap structure on FS. The result shows that the nodeless gap of $\beta$ pocket is related to the interlayer pairing within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals. We also find the formation of the inplane inter-orbital $d$-wave pairing between $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals that can enhance the dominated interlayer $s_{\pm}$-wave. This coexistence is robust symmetrically for $\mathrm{La_{3}Ni_{2}O_{7}}$ lattice.
@@
We thank the useful discussion with Wéi Wú, Guan-Hao Feng. This project was supported by NSFC-12494591, NSFC-92165204, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Guangdong Fundamental Research Center for Magnetoelectric Physics (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).
Supplementary Figure 1: Pairing configuration of $\mathrm{La_{3}Ni_{2}O_{7}}$ thin films.
Supplementary Figure 2: Superconducting gap $g_{t}^{xz}\Delta_{||}^{xz}$ projected onto FS.
To proceed with, we perform a transform to the bonding-antibonding representation as
$H_{\pm}=H_{||}\pm H_{\bot},\ \Delta_{\pm}=\Delta_{||}\pm\Delta_{\bot}.$ (23)
It is clear from this representation that, when along the Brillouin zone diagonal direction, the gaps for $\alpha,\beta,\gamma$ pockets respectively acquire the form
$\Delta_{||}^{x}(\cos k_{x}+\cos k_{y})+\Delta_{\bot}^{x},$ (24)
$\Delta_{||}^{x}(\cos k_{x}+\cos k_{y})-\Delta_{\bot}^{x},$ (25)
$-\Delta_{||}^{z}(\cos k_{x}+\cos k_{y})+\Delta_{\bot}^{z}.$ (26)
Eq. (26) shows that, in order to maximize the gap at $\gamma$ pocket around M=($\pi$,$\pi$) point, $-\Delta_{||}^{z}$, $\Delta_{\bot}^{z}$ have to be in antiphase. Such antiphase is also identified in Ref. [31], but there only $d_{x^{2}-y^{2}}$ orbital is relevant while $d_{z^{2}}$ orbital has been integrated out a priori. But for $d_{x^{2}-y^{2}}$ orbital in our case, as demonstrated in Eq. (24), they actually hold the same sign in order to maximize the gap at $\alpha$ pocket. For $\beta$ pocket in Eq. (25), a competition between $\Delta_{||}^{x},\Delta_{\bot}^{x}$ occurs, and our result as well as the experimental evidence suggest that $2\Delta_{||}^{x}\cos k_{x}<\Delta_{\bot}^{x}$. We also note that in Ref. [18], a negative sign of $\Delta_{||}^{x}$ is revealed using the functional renormalization group calculation for a multiorbital atomic Coulomb interaction. Since the pairing configuration in supplementary Fig. 1 mainly reflects the pairing preference for the $\mathrm{La_{3}Ni_{2}O_{7}}$ FS in the presence of $J_{\bot}$, the slight discrepancy may be attributed to other high-order processes besides $J_{\bot}$.